OpenStudy (anonymous):

Things just got complicated.. -1/2x + 2 = −x + 7

OpenStudy (rsadhvika):

\(\large \frac{-1}{2x+2} = -x+7\)

OpenStudy (rsadhvika):

its like this ?

OpenStudy (anonymous):

no negative 1 over 2 x + 2 = −x + 7

OpenStudy (rsadhvika):

\(\large \frac{-1}{2}x+2 = -x+7 \)

OpenStudy (rsadhvika):

like that ?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

i would start with multiplying the whole equation with -1 as i dont like negatives

OpenStudy (rsadhvika):

multiply the whole equation with -1

OpenStudy (rsadhvika):

\(\large \frac{1}{2}x-2 = x-7 \)

OpenStudy (rsadhvika):

subtract \(\frac{1}{2}x \) both sides

OpenStudy (anonymous):

how ?

OpenStudy (rsadhvika):

\(\large \frac{1}{2}x-2 = x-7 \) -1/2x -1/2x \(\large -2 = \frac{1}{2}x-7 \) add 7 both sides \(\large 5 = \frac{1}{2}x \) multiply 2 both sides 10 = x x = 10

OpenStudy (anonymous):

ohhhhhhh thanks